OpenStudy (anonymous):

The inverse of 5x+6 is (x-6)/5 right?

OpenStudy (wolf1728):

No - the inverse of 8 is 1/8

OpenStudy (freckles):

\[\text{ does } \frac{(5x+6)-6}{5} \text{ give you } x? \\ \text{ or } \\ 5(\frac{x-6}{5})+6 \text{ give you x}\]

OpenStudy (anonymous):

So im guessing im not right

OpenStudy (wolf1728):

I'd say you are incorrect

OpenStudy (freckles):

@wolf1728 is thinking about multiplicative inverses

OpenStudy (anonymous):

I know

OpenStudy (freckles):

and I think you mean the inverse of a function

OpenStudy (wolf1728):

Yes I am freckles

OpenStudy (freckles):

evaluate both of those expressions above

OpenStudy (freckles):

if you get x you are right

OpenStudy (freckles):

if you don't get x, then it would be incorrect

OpenStudy (anonymous):

I did: y = 5x + 6 x = 5y + 6 x - 6 = 5y (x - 6)/5 = y

OpenStudy (freckles):

your inverse is right and you can verify it by pluggin the inverse into the orginial and the orginial into the inverse as I did above and they should both evaluate to x

OpenStudy (anonymous):

And the second option you gave is correct

OpenStudy (anonymous):

Thanks for your help.

OpenStudy (freckles):

\[\text{ does } \frac{(5x+6)-6}{5} \text{ give you } x? \\ \text{ or } \\ 5(\frac{x-6}{5})+6 \text{ give you x} \\ \frac{(5x+6)-6}{5}=\frac{5x}{5}=x \\ 5(\frac{x-6}{5})+6=x-6+6=x\] the verification process proved you were right

OpenStudy (anonymous):

Oh! So like f(g(x)) and g(f(x)) with the inverse function and the function.

OpenStudy (freckles):

yes

OpenStudy (freckles):

if f and g are inverses then f(g(x))=x and g(f(x))=x